# Scaling 4D Representations

### v1

@@ 3.3 Baselines
Figure 7: The performance of the frozen 4DS-j model when attaching attention-based readouts at different layers of the model (as a percentage of the total number of self-attention blocks, 64 for 4DS-j as given in Tab. 1). The layer at 95% depth (60 for 4DS-j) offers a good compromise across tasks – we report results using this readout layer throughout the paper except for the SSv2 classification frozen evaluation where we use layer 75%.
@@ 4.1 Assessing video models 4D representations
Results for SSv2 action classification are reported by Top-1 accuracy as percentage; RE10k camera pose estimation uses Mean End Point Error (explained in supp. material, lower is better and minimum value is 0); Perception test point tracking uses Average Jaccard (higher is better, range 0-1); ScanNet depth prediction uses Absolute Relative Error (lower is better, minimum value is 0) and Waymo Open object tracking uses mean IoU (higher is better, range 0-1). We report accuracy and Average Jaccard in tables as percentage for better readability.
@@
Table 2 shows the quantitative results of the selected models – image models in first three rows, then video models from the literature and finally our models at the bottom.
@@
Image backbones. An interesting initial finding was that image backbones perform poorly regardless of model size on all tasks with the exception of depth estimation on ScanNet – likely because monocular cues dominate in short clips just over 1s long, where parallax is naturally limited and also where there are no moving objects. Dino does better than SigLIP across all tasks, hinting that CLIP-style representations may not be a good fit for 4D tasks. Finally both models are terrible at camera pose estimation.
@@
Video backbones. Text-supervised VideoPrism models show strong performance on SSv2 action classification and depth estimation, but are not as good on the tracking tasks and are quite poor on camera pose estimation. V-JEPA models shine on SSv2 and exhibit overall strong performance. Compared to the smaller v1, VideoMAE-v2 improves classification performance but at the expense of all other tasks.
4DS backbones. Our scaled up models, 4DS-G, e and j – explained in Sec. 3.1 – significantly improve over all the baselines, achieving top results on all tasks except SSv2 (only the 22B model 4DS-j is close to V-JEPA-H). The advantage is particularly noticeable when taking all tasks together as shown in Fig. 6.
Which layer is best for each task? As mentioned in the technical section, for 4DS we did not force a clear separation between encoder and decoder (except for a patchwise linear decoder). We investigated where in the model are the best representations for each task by trying a subset of 6 readout layers for each 4DS model at different model depths – 25%, 50%, 75%, 85%, 95% and 100% (e.g. 75% corresponds to layer 42 out of 56 for 4DS-e). Fig. 7 shows that for most tasks the layer at 95% depth seems like a good option. For the higher-level SSv2 task we used the layer at 75% of model depth for frozen evaluation; for finetuning this should not matter as the backbone can adapt so we used always 95%. These are the readout layers we used throughout the paper to report results.
@@
Finetuning evaluation. We repeated the evaluation using end-to-end finetuning – exactly the same as before but training both readout and backbone (instead of freezing the backbone). Note that this is significantly more expensive and also that we did not perform any test-time augmentation nor additional train-time augmentations compared to the frozen evaluation.
Results for a short finetuning schedule of 20k steps are shown in Tab. 3 and are consistent with the frozen evaluation results. There are big improvements over frozen evaluation in SSv2 and ScanNet, on others tasks results are comparable. Results with longer schedules are reported in Tab. 7 in the supplementary material.
Frozen and finetuning performance across 4DS model sizes is summarized visually in Fig. 8. In all cases performance continues to improve with model size.
@@ 4.2 Qualitative results
4DS-S (20M) 4DS-j (22B)
To understand the impact of model size, we visualize the outputs of our models during MAE training in Fig. 4(a) and across several downstream tasks, including depth estimation in Fig. 4(b), object tracking in Fig. 4(c), camera pose estimation in Fig. 5, and point tracking in Fig. 9 – we show results for all evals using a frozen backbone.
Across all tasks, larger models consistently demonstrate superior performance. During MAE training, larger models exhibit improved reconstruction quality. For example, in Fig. 4(a), the reconstruction of the spatula becomes progressively sharper as model size increases, suggesting that larger models learn more effective representations.
This trend holds true for downstream tasks as well. In depth estimation, smaller models produce washed out, inaccurate results. As model size increases, details emerge, such as the small patch of ground next to the kitchen cupboard in the second row and the patch of window glass not blocked by the curtain in the last row of Fig. 4(b). Camera pose estimation, a task requiring complex geometric reasoning, also benefits from increased model size – the task as is illustrated in Fig. 5 and here also larger models tend to be more accurate. Similarly, in point tracking, larger models significantly reduce errors, visualized as displacement errors, particularly around hands and moving objects in Fig. 9.
@@ 5 Conclusions
We have re-evaluated several state-of-the-art image and video models, trained with various types of supervision including language, on a set of mostly semantic-free tasks requiring strong 4D representations. We found that many of the improvements found in the respective papers do not translate well to these more geometrical / temporal scene understanding tasks and simple MAE from video still does quite well.
We then demonstrate that scaling MAE beyond what has been done in the literature ($\sim$1B parameters), all the way to 22B parameter models brings consistent improvement – we have a new collection of model checkpoints we call 4DS which may be useful for the community. This may open new avenues for further progress in vision similar to those that have been explored successfully in language models research.
Limitations. In this paper we used the most vanilla version possible of MAE – many improvements have been reported in the literature at small scale. As future work it would be interesting to identify which of these variations are most practical and fruitful when scaling, or if other learning approaches are superior. Another limitation of the paper is that we did not produce ”scaling laws” – but arguably provided evidence that this may soon be warranted.
@@
This supplementary material provides additional details on pretraining, evaluation protocol, tasks, readouts and baselines. It also includes additional qualitative results about SSv2, ScanNet and Waymo Open evaluations as well as extra MAE pretraining visualizations and a small study on correlation between the 5 tasks, when taking into account results from all baselines and 4DS models. Finally, we have a table with long end-to-end finetuning results for all models.
@@ A.2 Evaluation
A.2.1 Something-Something v2 action classification
@@
Readout details. The cross-attention readout module uses 768 parameters with 12 heads and a single learned query to predict logits for 174 classes. In training, we resize the shorter size of the video to 239 and take random temporal crop of shape 224x224 from it. We use colour augmentation with 0.8 probability of randomly adjusting the brightness, saturation, contrast and hue (see Tab. 6), and a 0.1 probability of converting to grayscale. In test time we take one 224x224 central crop from the video without any colour augmentation.
@@
A.2.2 RealEstate10k relative pose estimation
@@
Current SOTA. This is a new combination of dataset / task / metric we created as part of the research in this paper.
A.2.3 Perception Test point tracking
@@
A.2.4 Waymo Open object tracking
@@
A.2.5 ScanNet depth estimation
@@
Current SOTA. Some of the most recent methods include ZoeDepth [10] and Depth Anything [77]. Neither evaluates the method on ScanNet videos and they all evaluate zero-shot, which makes it hard to compare our work with these methods apples-to-apples. Nevertheless a strong recent method gets 0.093 AbsRel zero shot [74]. Another earlier work that trains and evaluates on ScanNet videos is Atlas which takes known camera intrinsics and extrinsics and achieves 0.061 AbsRel. Our best numbers with frozen features are 0.085 and with full finetuning 0.057 (we do not use camera parameters).
@@ A.3.1 SigLIP-2B
Architecture. This model consists of 48 self-attention pre-norm transformer layers, with 1536 feature dimension, and 16 attention heads. The transformer feed-forward block has a hidden dimension of 8192. CLS token?
@@ Appendix B Additional qualitative results
Figure 10: Top-1 accuracy of the 20M parameter 4DS-S (blue) vs the 4B parameter 4DS-e model (orange) on the SSv2 dataset and the DinoV2-g image model vs the 4DS-e model. Classes are ordered by the accuracy of the model shown in blue.
@@
Task correlation. The correlation coefficient for all evaluation tasks are provided in Fig. 14. ScanNet is the most unique, least correlated task – it is also the one where image backbones do best.
Figure 14: Pearson’s correlation coefficient for all frozen attention-based evaluation tasks across all models. This visualization is based on the results provided by Tab. 2. The RE10k and ScanNet values have been inverted, so that larger values correspond to better performance, consequently all correlations are positive.
Appendix C Finetuning Results
@@
Model Size (M) Image (I)/ Video (V) SSv2 RE10k Percep. test ($\uparrow$) ScanNet Waymo Open ($\uparrow$)
($\uparrow$) ($\downarrow$) ($\downarrow$)
SigLIP [15] 1,705 I 58.9 / 57.8 2.58 / 2.26 37.2 / 42.5 1.07 / 0.94 49.6 / 52.6
DinoV2-L [50] 303 I 62.3 / 62.0 1.96 / 2.05 53.9 / 56.0 0.8 / 0.75 56.0 / 60.3
DinoV2-g [50] 1,135 I 63.9 / 62.3 2.00 / 2.06 47.5 / 50.4 0.69 / 0.66 55.5 / 59.8
VideoPrism-B [80] 114 V 64.4 / 64.3 0.70 / 0.68 76.1 / 77.4 1.01 / 0.95 71.3 / 74.2
VideoPrism-g [80] 1,113 V 68.9 / 68.3 0.75 / 0.73 77.6 / 78.5 0.83 / 0.79 71.8 / 74.3
V-JEPA-L [8] 307 V 71.5 / 71.1 0.28 / 0.26 80.7 / 82.4 0.88 / 0.82 78.3 / 79.9
V-JEPA-H [8] 635 V 73.4 / 72.7 0.28 / 0.27 78.0 / 81.4 0.8 / 0.76 78.8 / 80.9
VideoMAE-B [66] 87 V 57.2 / 58.0 0.33 / 0.30 82.1 / 83.6 1.35 / 1.24 76.4 / 78.2
VideoMAE-L [66] 305 V 66.7 / 66.8 0.27 / 0.28 81.5 / 82.3 0.94 / 0.85 78.7 / 80.9
VideoMAE-H [66] 633 V 67.4 / 67.5 0.30 / 0.29 79.7 / 81.1 0.86 / 0.81 77.5 / 79.8
VideoMAEv2-g [70] 1,013 V 70.5 / 69.9 0.39 / 0.38 80.4 / 82.0 0.88 / 0.81 77.0 / 80.0
4DS-S 24 V 43.9 / 46.0 0.59 / 0.51 77.4 / 78.3 1.88 / 1.74 73.5 / 75.7
4DS-B 91 V 54.9 / 55.4 0.43 / 0.41 81.7 / 82.9 1.39 / 1.3 76.3 / 78.5
4DS-L 310 V 63.8 / 64.3 0.33 / 0.30 83.9 / 85.2 0.99 / 0.94 79.2 / 81.1
4DS-H 639 V 66.4 / 66.6 0.31 / 0.29 84.6 / 85.6 0.88 / 0.84 80.1 / 82.0
4DS-G 1,848 V 68.5 / 68.5 0.25 / 0.24 85.1 / 85.9 0.76 / 0.72 81.0 / 82.4
4DS-e 3,811 V 69.6 / 69.0 0.24 / 0.24 85.1 / 85.8 0.69 / 0.66 81.5 / 82.8
4DS-j 21,495 V 72.5 / 72.0 0.25 / 0.26 86.3 / 86.6 0.59 / 0.57 81.3 / 83.0

### v2

@@ 3.4 Baselines
Figure 6: The performance of the frozen 4DS-j model when attaching attention-based readouts at different layers of the model (as a percentage of the total number of self-attention blocks, 64 for 4DS-j as given in Tab. 1). The layer at 95% depth (60 for 4DS-j) offers a good compromise across tasks – we report results using this readout layer throughout the paper except for action classification frozen evaluation where we use layer 75%.
@@ 4.1 Assessing video models 4D representations
RE10k camera pose estimation uses Mean End Point Error (explained in supp. material, lower is better and minimum value is 0); Perception test point tracking uses Average Jaccard (higher is better, range 0-1); ScanNet depth prediction uses Absolute Relative Error (lower is better, minimum value is 0) and Waymo Open object tracking uses mean IoU (higher is better, range 0-1). Results for SSv2 and Kinetics700-2020 action classification are reported by Top-1. We report accuracy and Average Jaccard in tables as percentages for better readability.
@@
Table 2 shows the quantitative results of the selected models: image models in first three rows, then existing video models, and finally our 4DS models at the bottom.
@@
Image backbones. An interesting initial finding was that image backbones perform poorly regardless of model size on all 4D tasks with the exception of depth estimation on ScanNet – likely because monocular cues dominate in short clips just over 1s long, where parallax is naturally limited and also where there are no moving objects. Dino does less poorly than SigLIP across all tasks, hinting that CLIP-style representations may not be a good fit for 4D tasks. All models are terrible at camera pose estimation, but they are good at Kinetics, which is known to depend strongly on the quality of high-level (semantic) appearance features.
@@
Video backbones. Text-supervised VideoPrism models show strong performance on depth estimation, but are not as good on the tracking tasks and are quite poor on camera pose estimation. V-JEPA models exhibit overall strong performance. Compared to the smaller v1, VideoMAE-v2 uses additional language-pretraining which seems to lead to worse performance on the 4D tasks. Language pretrained models do however very well on the semantic tasks, SSv2 and Kinetics700-2020. VJEPA does great on SSv2.
4DS backbones. Our scaled up models, 4DS-G, e and j – explained in Sec. 3.1 – significantly improve over all the baselines, achieving top results on all 4D tasks. On semantic tasks they are not as competitive (although results continuously improve with scale), illustrating why prior self-supervised video learning work, which has focused on those tasks, has not pursued scaling. The smaller 4DS models are not as good as corresponding VideoMAE models. Partly this may be because 4DS model features are loaded from the layer at 95% depth, so from slightly shallower features and without an extra decoder, as well as the other simplifications we introduced.
Distilling from large models. We tried distilling a 4DS-e (4B parameters) into a B model (90M parameters). Results are shown in Tab. 2 and suggest that one can also obtain SOTA small models by distilling from the larger ones. In detail – we input the same video to both the randomly initialized 4DS-B model and the pretrained 4DS-e frozen model, without dropping any patches. We trained the 12 layer B model to approximate the activations of two different layers of the teacher model, for 100k steps using batch size 1024 and the L1 loss. We added two parallel linear layers at the end of the B model: one predicted teacher layer 36 activations, the other one layer 51 activations (note 4DS-e has many more channels than B, so adapter layers were unavoidable); afterwards we discarded these extra linear layers – they were only used during distillation. The improvement over 4DS-B model trained by itself is quite noticeable and raises interesting questions for future work.
@@
Finetuning evaluation. We repeated the evaluation using end-to-end finetuning – exactly the same as before but training both readout and backbone (instead of freezing the backbone). Note that this is significantly more expensive and also that we did not perform any test-time augmentation nor additional train-time augmentations relative to the frozen evaluation. Results for a short finetuning schedule of 20k steps are shown in Tab. 3 and are consistent with the frozen evaluation results. There are big improvements over frozen evaluation in ScanNet and SSv2, on others tasks results are comparable. Results with longer schedules are reported in Tab. 9 in the supplementary material. Frozen and finetuning performance across 4DS model sizes is summarized visually in Fig. 7. In all cases performance continues to improve with model size.
4.2 Other studies
Which layer is best for each task? As mentioned in the technical section, for 4DS we did not force a clear separation between encoder and decoder (except for a patchwise linear decoder). We investigated where in the model are the best representations for each task by trying a subset of 6 readout layers for each 4DS model at different model depths – 25%, 50%, 75%, 85%, 95% and 100% (e.g. 75% corresponds to layer 42 out of 56 for 4DS-e). Fig. 6 shows that for most tasks the layer at 95% depth seems like a good option. For the semantic tasks (SSv2, Kinetics), we used the layer at 75% of model depth for frozen evaluation; for finetuning we also used 95% except for Kinetics where we used 75%. These are the readout layers we used throughout the paper to report results.
@@ 4.2 Other studies
Data scaling impact. We also evaluated the value of the amount of training. Results for training over 250M, 500M and 1B examples for model 4DS-e is shown in Tab. 4. There are nearly always gains from extra training.
Decoding hyperparameters. We include an ablation confirming our defaults in the supp. material, sec. C.
# examples 250M 500M 1B
RE10k 0.35 0.29 0.27
Perception Test 81.8 82.9 82.4
ScanNet 1.39 1.11 0.94
Waymo Open 74.7 76.8 78.0
SSv2 59.0 64.3 68.2
Table 4: Impact of number of examples seen in pretraining on downstream performance for an 4DS-e model (4B parameters).
@@ 4.3 Qualitative results
To further help understand the impact of model size, we visualize the outputs of our models during MAE training in Fig. 4(a) and on depth estimation in Fig. 4(b), object tracking in Fig. 4(c) and camera pose estimation in Fig. 5. All obtained using a frozen backbone.
Larger models consistently demonstrate superior performance. During MAE training, they exhibit improved reconstruction quality: in Fig. 4(a), the reconstruction of the spatula becomes progressively sharper as model size increases. This trend holds true for downstream tasks as well. In depth estimation, smaller models produce washed out, inaccurate results. As model size increases, details emerge (see Fig. 4(b)). Camera pose estimation results are illustrated in Fig. 5 (larger models also tend to be more accurate).
@@ 5 Conclusions
We have re-evaluated several state-of-the-art image and video models, trained with various types of supervision including language, on a set of mostly semantic-free tasks requiring strong 4D representations. We found that many of the improvements found in the respective papers do not translate well to these more geometrical / temporal scene understanding tasks.
We then demonstrate that scaling MAE beyond what has been done in the literature ($\sim$1B parameters), all the way to 22B parameter models brings consistent improvement – we have a new collection of model checkpoints we call 4DS which may be useful for the community. The largest models may not be practical for all applications, but it is likely they can be distilled to smaller form factors while preserving much of their performance. Generally, we hope that the scaling insights in this paper may open new avenues for further progress in vision similar to those that have been explored successfully in language models research.
Limitations. In this paper we used the most vanilla version possible of MAE (which we called SimpleMAE) – many improvements have been reported in the literature at small scale. As future work it would be interesting to identify which of these variations are most practical and fruitful when scaling, or if other learning approaches are superior. Another limitation of the paper is that we did not produce ”scaling laws” – but arguably provided evidence that this may soon be warranted.
@@
This supplementary material provides additional details on pretraining, evaluation protocol, tasks, readouts and baselines. It also includes additional qualitative results about SSv2, ScanNet and Waymo Open evaluations as well as extra MAE pretraining visualizations and a small study on correlation between the 5 tasks, when taking into account results from all baselines and 4DS models. Finally, we have an ablation study on decoding hyperparameters and a table with long end-to-end finetuning results for all models.
@@ A.2 Evaluation
K700-2020 Classification
LearnedQueries(num_channels=1024)
CrossAttention(qkv_size=1024, num_heads=16)
Linear(output_size=700)
12,281,532
@@
A.2.1 Action classification 1: Something-Something v2
@@
Readout details. The cross-attention readout module uses 768 channels with 12 heads and a single learned query to predict logits for 174 classes. In training, we resize the shorter size of the video to 239 and take random temporal crop of shape 224x224 from it. We use colour augmentation with 0.8 probability of randomly adjusting the brightness, saturation, contrast and hue (see Tab. 7), and a 0.1 probability of converting to grayscale. In test time we take one 224x224 central crop from the video without any colour augmentation.
@@
A.2.2 Action classification 2: Kinetics-700-2020
Task definition. Given a video clip of 16 frames of resolution 224x224 with stride 2, the model is tasked to predict one action class out of 700. At test time, since videos are variable length and up to 10s long (250 total frames at 25 fps), we first loop the video as many times as necessary to get 250 frames, then take 7 linearly spaced 16-frame clips and pass them through the model. Finally, we average the softmax scores across all of them. Top-1 accuracy is used to measure the performance.
Readout details. The same readout is used as for SSv2, but slightly larger – 1024 channels and 16 heads instead of 768 and 12.
Current SOTA. Vision-language models are very strong in Kinetics. Among them, VideoPrism [82] is one of the strongest models for frozen attention-based evaluation on kinetics-400 (an earlier, smaller version of the Kinetics-700-2020 version we use). It is also the strongest one in our evaluation.
A.2.3 Relative pose estimation: RealEstate10k
@@
Current SOTA. This is a new combination of dataset / task / metric we created as part of the research in this paper. As a strong baseline, zero shot DUSt3R [73] gets EPE of 1.06, which is worse than our model (0.24). This is because Dust3r sometimes fails to register the two pointmaps producing large errors.
A.2.4 Point tracking: Perception Test
@@
A.2.5 Object tracking: Waymo Open
@@
A.2.6 Depth estimation: ScanNet
@@
Current SOTA. Some of the most recent methods include ZoeDepth [10] and Depth Anything [79]. Neither evaluates the method on ScanNet videos and they all evaluate zero-shot, which makes it hard to compare our work with these methods apples-to-apples. Among these, DUSt3R [73] zero-shot gets 0.088 whereas a slightly older method [76] gets 0.093. An earlier work that trains and evaluates on ScanNet videos is Atlas [49] which takes known camera intrinsics and extrinsics and achieves 0.061 AbsRel. Our best numbers with frozen features are 0.084 and with full finetuning 0.057 (we do not use camera parameters).
@@ A.3.1 SigLIP-2B
Architecture. This model consists of 48 self-attention pre-norm transformer layers, with 1536 feature dimension, and 16 attention heads. The transformer feed-forward block has a hidden dimension of 8192.
@@ Appendix B Additional qualitative results
Figure 8: Top-1 accuracy of the 20M parameter 4DS-S (blue) vs the 4B parameter 4DS-e model (orange) on the SSv2 (above) and Kinetics700-2020 (below) datasets and the DinoV2-g image model vs the 4DS-e model. Classes are ordered by the accuracy of the model shown in blue.
@@
Perception Test point tracking. Additional qualitative results for Perception Test point tracking are provided in
4DS-S (20M) 4DS-j (22B)
. Larger models significantly reduce errors, visualized as displacement errors, particularly around hands and moving objects.
@@
Task correlation. The correlation coefficient for all evaluation tasks are provided in Fig. 14. There is a positive correlation with model size suggesting that larger model size leads to better performance. Kinetics classification is the least correlated task.
Additional relative performance visualization. A radar plot showing relative performance visually between some of the strongest models in our comparison is shown in Fig. 13.
Figure 14: Pearson’s correlation coefficients for all frozen attention-based evaluation tasks across all models. This visualization is based on the results provided by Tab. 2. The RE10k and ScanNet values have been inverted, so that larger values correspond to better performance. A highly semantic task, K700-2020, displays negative correlation with most other non-semantic tasks. Model size has a positive correlation with all tasks suggesting that larger model size leads to better performance.
Appendix C Decoding hyperparameters study
There are two main hyperparameters in our approach to decoding in SimpleMAE: the decoder patch size and the number of latent layers (where latent tokens self-attend with regular tokens). Results for 4DS-e are shown in 8 and validate the default values we used for all models but 4DS-j – which due to memory constraints uses 2 latent layers and patch size (4,32,32). Presumably with larger accelerators or a more efficient implementation it would be possible to get extra performance out of this model using the same hyperparameters as for all other models.
Table 8: Study on decoding hyperparameters using the RE10k, ScanNet and SSv2 tasks, pre-training the 4DS-e model on 1B examples (with quick evaluation on 4096 test examples for each task). 4 latent layers and (2,16,16) patches seems preferable in all cases.
Task Decoding patch size # latent layers
2 4
RE10k (2,16,16) 0.32 0.29
(4,32,32) 0.32 0.33
ScanNet (2,16,16) 1.0 0.95
(4,32,32) 1.15 1.12
SSv2 (2,16,16) 63.3 65.8
(4,32,32) 60.6 61.9
Appendix D Extra Finetuning Results
@@
Model Size (M) Image (I)/ Video (V) SSv2 K700-2020 RE10k Percep. test ($\uparrow$) ScanNet Waymo Open ($\uparrow$)
($\uparrow$) ($\uparrow$) ($\downarrow$) ($\downarrow$)
SigLIP [15] 1,705 I 58.9 / 57.8 68.7 / 68.5 2.58 / 2.26 37.2 / 42.5 1.14 / 1.02 49.6 / 52.6
DinoV2-L [50] 303 I 62.3 / 62.0 64.6 / 67.6 1.96 / 2.05 53.9 / 56.0 0.77 / 0.74 56.0 / 60.3
DinoV2-g [50] 1,135 I 63.9 / 62.3 67.0 / 68.8 2.00 / 2.06 47.5 / 50.4 0.67 / 0.66 55.5 / 59.8
VideoPrism-B [82] 114 V 64.4 / 64.3 61.3 / 63.9 0.70 / 0.68 76.1 / 77.4 0.99 / 0.94 71.3 / 74.2
VideoPrism-g [82] 1,113 V 68.9 / 68.3 71.7 / 72.7 0.75 / 0.73 77.6 / 78.5 0.81 / 0.78 71.8 / 74.3
V-JEPA-L [8] 307 V 71.5 / 71.1 60.3 / 62.6 0.28 / 0.26 80.7 / 82.4 0.87 / 0.82 78.3 / 79.9
V-JEPA-H [8] 635 V 73.4 / 72.7 61.5 / 63.7 0.28 / 0.27 78.0 / 81.4 0.79 / 0.76 78.8 / 80.9
VideoMAE-B [67] 87 V 57.2 / 58.0 41.2 / 45.0 0.33 / 0.30 82.1 / 83.6 1.34 / 1.24 76.4 / 78.2
VideoMAE-L [67] 305 V 66.7 / 66.8 53.8 / 57.0 0.27 / 0.28 81.5 / 82.3 0.91 / 0.85 78.7 / 80.9
VideoMAE-H [67] 633 V 67.4 / 67.5 56.8 / 58.2 0.30 / 0.29 79.7 / 81.1 0.85 / 0.79 77.5 / 79.8
VideoMAEv2-g [71] 1,013 V 70.5 / 69.9 70.5 / 70.3 0.39 / 0.38 80.4 / 82.0 0.84 / 0.79 77.0 / 80.0
4DS-S 24 V 43.9 / 46.0 30.8 / 33.9 0.59 / 0.51 77.4 / 78.3 1.89 / 1.72 73.5 / 75.7
4DS-B 91 V 54.9 / 55.4 38.6 / 42.8 0.43 / 0.41 81.7 / 82.9 1.38 / 1.30 76.3 / 78.5
4DS-L 310 V 63.8 / 64.3 48.7 / 52.1 0.33 / 0.30 83.9 / 85.2 0.99 / 0.93 79.2 / 81.1
4DS-H 639 V 66.4 / 66.6 51.9 / 54.4 0.31 / 0.29 84.6 / 85.6 0.89 / 0.83 80.1 / 82.0
4DS-G 1,848 V 68.5 / 68.5 55.3 / 57.8 0.25 / 0.24 85.1 / 85.9 0.75 / 0.70 81.0 / 82.4
4DS-e 3,811 V 69.6 / 69.0 55.2 / 60.4 0.24 / 0.24 85.1 / 85.8 0.68 / 0.65 81.5 / 82.8
4DS-j 21,495 V 72.5 / 72.0 59.2 / 62.9 0.25 / 0.26 86.3 / 86.6 0.59 / 0.57 81.3 / 83.0